Two adjacent sides of a parallelogram are 24 cm and 18cm . If the distance between longer side is 12 cm , find the distance between shorter sides.

Grade:8

1 Answers

Arun
6 years ago
 
Given : Two adjacent sides of the parallelogram are AB = 24 cm and BC = 18 cm and the distance between the longer side, DE = 12 cm.
To find : The distance between the shorter side.
Let the distance between the shorter side, DF be 'x' cm.
Area of the parallelogram = Base × Height 
Area of the parallelogram, considering AB as the base = AB × DE
= 24 × 12
Area = 288 cm²
Now considering BC as the base, the area of the parallelogram = BC × DF
= 18 × x
Area = 18x
Since the area of the parallelogram will be the same, if we consider Ab as the base or BC as the base.
So, AB × DE = BC × DF
⇒ 288 18x
x = 288/18
x = 16 cm
So, the distance between the shorter side is 16 cm